# Calculation of harmonic oscillator brackets in SU(3) basis

### v1

@@ 2 Definitions and motivation: HOB in SO(3) basis
in [5]. Here $d$ is a nonnegative real number. The variables in the brackets are arranged in a fixed way and the orthogonal transformation of the coordinates $\boldsymbol{r}_{1},\boldsymbol{r}_{2}$ is written as the two-dimensional matrix
@@
This two-dimensional matrix is called the Talmi-Moshinsky transformation. It is parametrized by the variable $d.$ In [5] the original expression for HOBs derived by B.Buck and A.C.Merchant [16]
@@
An additional component in the eq.(14) is the 9j symbol, denoted by the curly brackets. The 9j symbol in the expression has eight summation indices, although only three are independent. The described expression has proven to be useful in the ab initio nuclear structure calculations [17, 18, 19, 20, 21]. It is well-balanced and gives fast and precise results. However, the calculation of the 9j coefficient is still expensive, so an alternative expression that could producen nice results would be useful.
@@ 4 The Talmi-Moshinsky transformation in $SU(3)$ representation
d+1d 1+1d
1+1d -d+1d
=(1001)×(d1+d11+d11+dd1+d)
d+1d 1+1d
1+1d -d+1d
[…]
@@
cos(θ) sin(θ)
-sin(θ) cos(θ)
cos(θ) sin(θ)
-sin(θ) cos(θ)
[…]
@@
The action of the operator $\hat{a}$ results in a Wigner d-matrix of the irreducible representation of the group $SU(2)$, which is a subgroup of $SU(3)$. It is an orthogonal transformation that transforms a function of the coupled $SU(3)$ irreps $e_{1}$ and $e_{2}$ depending on the variables $\boldsymbol{r}_{1}$ and $\boldsymbol{r}_{2}$ into a function depending on the variables $\boldsymbol{r}_{1}^{\prime}$ and $\boldsymbol{r}_{2}^{\prime}$. The Wigner d-matrix depends on the same angle $\theta$. We change the notation a little bit, folowing the eq.(25), namely
@@
We rearange this expression to make it computationally more feasible by using $\Gamma$ function instead of factorials and pulling out terms of equation independent from the summation
@@
from [26], which we also rewrite in the $E\alpha J$ notation as
@@ 5 The calculation of the HOB
In this section, we gather the presented knowledge into a new approach for the calculation of the HOB. HOB is defined as $\langle e_{1}l_{1},e_{2}l_{2}:L|e_{1}^{\prime}l_{1}^{\prime},e_{2}^{\prime}l_{2}^{\prime}:L\rangle_{d}$, or $\langle e_{1}l_{1},e_{2}l_{2}:L|T(d)|e_{1}^{\prime}l_{1}^{\prime},e_{2}^{\prime}l_{2}^{\prime}:L\rangle.$ We transform the $SO(3)$ basis into the $SU(3)$ basis by using the Clebsh-Gordan isofactors and summing though the intermediate states
@@
We are left with two summations of permitted $J$ and $\alpha$ values with bounds $|M|\ \leq J\leq E/2$ and $|M^{\prime}|\ \leq J\leq E/2$. The multiplicity label $\alpha$ enumerates the reccuring $SO(3)$ irrep $L$ in the $SU(3)$ irrep $J$. The CG coefficients in the SO(3) bases are constructed subsequently from the overlaps between the SU(2) and SO(3) bases.
@@ E $\delta{(SU3HOB)}$ $\delta{(HOB)}$
We see that for small HO E up until E=12, the HOB code is more accurate, however for larger E, the SU3HOB takes the lead. More detailed analysis of data shows that for larger matrices where L values are up until approximately $\frac{E}{2},$ the HOB code deteriotes faster. However, for larger Ls, the SU3HOB code gives worse results (see table (3)). This can be attributed to the SU3 Clebsh-Gordan code, which has trouble coupling for the larger Ls. We checked the errors for the Clebsh-Gordan matrix and the d-matrix used for matrix multiplication code following the same condition of orthonormality. The results are presented in table (4). We see a clear deterioration of the CG matrix, while the d matrix is pretty stable.
@@ E $\delta{(CGMat)}$ $\delta{(Djmm^{\prime})}$
Table 4: Error of the Clebsh-Gordan matrix compared to Wigner d-matrix for a given HO Energy quanta.

### v2

@@ 2 Definitions and motivation: HOB in SO(3) basis
in [5]. Here $d$ is a nonnegative real number, parameterizing the Talmi-Moshinsky transformation (see eq.(5)). The variables in the brackets are arranged in a fixed way and the orthogonal transformation of the coordinates $\boldsymbol{r}_{1},\boldsymbol{r}_{2}$ is written as the two-dimensional matrix
@@
This two-dimensional matrix is called the Talmi-Moshinsky transformation. In [5] the original expression for HOBs derived by B.Buck and A.C.Merchant [16]
@@
An additional component in the eq.(14) is the 9j symbol, denoted by the curly brackets. The 9j symbol in the expression has eight summation indices, although only three are independent. The described expression has proven to be useful in the ab initio nuclear structure calculations [17, 18, 19, 20, 21]. It is well-balanced and gives fast and precise results. However, the calculation of the 9j coefficient is still expensive, so an alternative expression that could produce nice results would be useful.
@@ 4 The Talmi-Moshinsky transformation in $SU(3)$ representation
($\sqrt{\frac{d}{1+d}}$&$\phantom{-}\sqrt{\frac{1}{1+d}}$\\ $\sqrt{\frac{1}{1+d}}$&$-\sqrt{\frac{d}{1+d}}$\\ \end{tabular}\right)=\left(\begin{tabular}[]{ll}$1$&$\phantom{-}0$\\ $0$&$-1$\\ \end{tabular}\right)\times\left(\begin{tabular}[]{ll}$\phantom{-}\sqrt{\frac{d}{1+d}}$&$\sqrt{\frac{1}{1+d}}$\\ $-\sqrt{\frac{1}{1+d}}$&$\sqrt{\frac{d}{1+d}}$\\ \end{tabular}\right).$ (29)
@@
($\phantom{-}\cos(\theta)$&$\sin(\theta)$\\ $-\sin(\theta)$&$\cos(\theta)$\\ \end{tabular}\right).$ (32)
@@
The action of the operator $\hat{a}$ results in a Wigner d-matrix of the irreducible representation of the group $SU(2)$, which is a subgroup of $SU(3)$. It is an orthogonal transformation that transforms a function of the coupled $SU(3)$ irreps $e_{1}$ and $e_{2}$ depending on the variables $\boldsymbol{r}_{1}$ and $\boldsymbol{r}_{2}$ into a function depending on the variables $\boldsymbol{r}_{1}^{\prime}$ and $\boldsymbol{r}_{2}^{\prime}$. The Wigner d-matrix depends on the same angle $\theta$. We change the notation a little bit, following the eq.(25), namely
@@
We rearrange this expression to make it computationally more feasible by using $\Gamma$ function instead of factorials and pulling out terms of equation independent from the summation
@@
from [26]. which we also rewrite in the $E\alpha J$ notation as
@@ 5 The calculation of the HOB
In this section, we gather the presented knowledge into a new approach for the calculation of the HOB. HOB is defined as $\langle e_{1}l_{1},e_{2}l_{2}:L|e_{1}^{\prime}l_{1}^{\prime},e_{2}^{\prime}l_{2}^{\prime}:L\rangle_{d}$, or $\langle e_{1}l_{1},e_{2}l_{2}:L|T(d)|e_{1}^{\prime}l_{1}^{\prime},e_{2}^{\prime}l_{2}^{\prime}:L\rangle.$ We transform the $SO(3)$ basis into the $SU(3)$ basis by using the Clebsch-Gordan isofactors and summing though the intermediate states
@@
We are left with two summations of permitted $J$ and $\alpha$ values with bounds $|M|\ \leq J\leq E/2$ and $|M^{\prime}|\ \leq J\leq E/2$. The multiplicity label $\alpha$ enumerates the reccuring $SO(3)$ irrep $L$ in the $SU(3)$ irrep $J$. The label $\alpha$ is the same as the label $K$ introduced by J. P. Elliott in [27]. The CG coefficients in the SO(3) bases are constructed subsequently from the overlaps between the SU(2) and SO(3) bases.
@@ E $\delta{(SU3HOB)}$ $\delta{(HOB)}$
We see that for small HO E up until E=12, the HOB code is more accurate, however for larger E, the SU3HOB takes the lead. More detailed analysis of data shows that for larger matrices where L values are up until approximately $\frac{E}{2},$ the HOB code deteriotes faster. However, for larger Ls, the SU3HOB code gives worse results (see table (3)). This can be attributed to the SU3 Clebsch-Gordan code, which has trouble coupling for the larger Ls. We checked the errors for the Clebsch-Gordan matrix and the d-matrix used for matrix multiplication code following the same condition of orthonormality. The results are presented in table (4). We see a clear deterioration of the CG matrix, while the d matrix is pretty stable.
@@ E $\delta{(CGMat)}$ $\delta{(Djmm^{\prime})}$
Table 4: Error of the Clebsch-Gordan matrix compared to Wigner d-matrix for a given HO Energy quanta.